In a bag there are 10 balls in which 3 are red. The probability that there will be atleast one red ball,in a draw of two balls is?

Grade:12

1 Answers

Arun
4 years ago
also want to know theanswer
Dear student
 
probability of drawing non red ball in 1st time = 7/10
 
probability of drawing non red ball in succession = 7/10 * 6/9 = 42/90
 
hence
 
probability of drawing atleast one red ball = 1 – 42 /90 = 48/90
 
I am not sure about my answer. 
Pleaselet me know if this is correct or incorrect.
Iw